OpenStudy (anonymous):

how to solve this identity (cosx)^4 -(Sin X)^4 = Cos(2x)

OpenStudy (anonymous):

factor

OpenStudy (anonymous):

\[(a^4-b^4)=(a^2+b^2)(a^2-b^2)\]

OpenStudy (anonymous):

Ok, I just factor that and simplify?

OpenStudy (anonymous):

yeah i used \(a\) for cosine and \(b\) for sine

OpenStudy (anonymous):

the first term \[a^2+b^2\] is the same as \[\cos^2(x)+\sin^2(x)\] which is 1

OpenStudy (anonymous):

so you really just get \[\cos^2(x)-\sin^2(x)\]

OpenStudy (anonymous):

I'm trying to see what you are telling me to do? a little confuse.